# PartSTAD: 2D-to-3D Part Segmentation Task Adaptation

### v1

@@ Abstract
We introduce PartSTAD, a method designed for the task adaptation of 2D-to-3D segmentation lifting. Recent studies have highlighted the advantages of utilizing 2D segmentation models to achieve high-quality 3D segmentation through few-shot adaptation. However, previous approaches have focused on adapting 2D segmentation models for domain shift to rendered images and synthetic text descriptions, rather than optimizing the model specifically for 3D segmentation. Our proposed task adaptation method finetunes a 2D bounding box prediction model with an objective function for 3D segmentation. We introduce weights for 2D bounding boxes for adaptive merging and learn the weights using a small additional neural network. Additionally, we incorporate SAM, a foreground segmentation model on a bounding box, to improve the boundaries of 2D segments and consequently those of 3D segmentation. Our experiments on the PartNet-Mobility dataset show significant improvements with our task adaptation approach, achieving a 7.0%p increase in mIoU and a 5.2%p improvement in mAP50 for semantic and instance segmentation compared to the SotA few-shot 3D segmentation model.
@@ 3.1 PartSLIP [32]
Voting for Super Points
@@ 4 PartSTAD: Task Adaptation for 2D-to-3D
In the appendix, we demonstrate that the choice of the loss function, as compared to other alternatives such as the cross-entropy loss, is crucial to achieving a substantial improvement in our task adaptation.
@@ 4.3 SAM [20] Mask Integration
Input
GT
SATR [2]
SATR [2]+SP
PartSLIP [32]
PartSTAD (Ours)
@@
Method
Mean
Storage Furniture
Table Faucet Chair Switch
Trash Can
Toilet
Eye- glasses
Globe Cart Window Bottle
Dis- penser
Laptop
Coffee Machine
USB Remote Pen
Dish- washer
Scissors
SATR [2]
29.3
20.6
23.3
37.1
33.1
21.4
11.6
17.6
63.5
83.1
30.8
67.7
50.0
38.2
11.2
20.0
30.2
17.2
51.1
12.9
36.8
SATR [2]+SP
34.8
28.9
28.0
39.8
37.7
37.0
15.9
22.1
74.2
83.5
37.2
69.2
55.5
41.4
12.4
23.2
33.4
28.0
57.3
18.8
43.0
PartSLIP [32]
58.0
52.3
44.6
66.2
82.8
52.1
23.3
50.4
89.2
92.9
78.7
78.3
80.7
70.8
31.2
37.2
52.1
36.6
68.9
59.1
61.4
@@
w/o Weight Pred.
61.9
56.6
45.0
65.5
85.0
51.9
24.3
56.6
88.7
98.1 85.4
72.6
85.9
70.5
31.5
36.6
57.1
36.6
64.3
61.4
60.8
w/o SAM Integration
62.1
54.0
45.7
64.0
83.1
53.0
18.4
49.4
92.1
95.7
81.2
78.6
81.0
72.0
33.6
37.1
53.6
46.8
65.1
59.2
67.5
PartSTAD (Ours)
65.0 59.5 47.8
65.3
85.3 57.9
21.2
57.5 92.5
93.5
85.0
76.1
83.6
73.7 34.6
35.8
59.9 53.4
60.6
60.5
68.5
Table 1: Quantitative comparison of the semantic segmentation results (mIoU(%)) on the PartNet-Mobility dataset. Ours achieves a 7.0%p improvement in average mIoU compared to PartSLIP [32], the SotA few-shot 3D segmentation method, while consistently increasing mIoU across all categories. Please refer to the appendix for the complete table with results for all 45 categories.
@@ 5 Experiment Results
In the experiments, we compare our method with the SotA 2D-prior-based zero-shot/few-shot 3D segmentation models using the PartNet-Mobility [55] dataset. In the appendix, we present additional results on the ablation study, outcomes with scanned data, and comprehensive results covering all PartNet-Mobility categories.
@@ 5.1 Experiment Setup
Dataset
We use the PartNet-Mobility [55] dataset, which is also used in the experiments of PartSLIP [32]. The PartNet-Mobility [55] dataset includes 45 object categories. For each category, the training split includes 8 shapes, while the test split has a range of the number of objects from 6 to 338, totaling 1,906 objects across all 45 categories. Following PartSLIP, 8 shapes from the test set of each category are designated as the validation set. To obtain the 2D bounding boxes for each object, we use 10 images from fixed viewpoints across all our experiments.
Evaluation Metrics
@@
[…]
Baselines
@@
Ablation Study
We also conduct an ablation study, comparing our method with two major components being ablated: the weight prediction (Sec. 4.2) and the SAM mask integration (Sec. 4.3). In the appendix, we also compare our method with cases using cross-entropy loss instead of mRIoU loss (Sec. 4.1) and using GLIP bounding box confidence score instead of our learned weights in the weighted voting (Sec. 4.2).
@@ 5.2 Semantic Segmentation Results
Tab. 1 and Fig. 3 present the quantitative and qualitative comparisons of part semantic segmentation results, respectively. Please refer to the appendix for the complete table encompassing all 45 categories. In comparison to PartSLIP [32], a SotA few-shot 3D segmentation method, we achieve a significant 7.0%p mIoU improvement over the entire set of 45 categories. Furthermore, for each specific category, we consistently demonstrate improvement, surpassing by more than 15%p in some categories (e.g., Remote). Qualitatively, the results also highlight the effectiveness of our task adaptation and the SAM mask integration in segmenting parts more precisely, even for small (e.g., Camera, Chair) and thin (e.g., Clock) parts of objects.
@@
Input
GT
SAM3D [62]
PartSLIP [32]
PartSTAD (Ours)
@@
Method
Mean
Storage Furniture
Table Faucet Chair
Trash Can
Switch Toilet
Eye- glasses
Globe Cart Window
Dis- penser
Bottle Laptop
Coffee Machine
USB Remote Pen
Dish- washer
Scissors
PartSLIP [32]
41.6
29.1
32.6
43.0
82.2
11.6
21.2
36.2
69.8
92.1
69.5
23.6
56.3
73.4
17.8
15.8
20.9
19.9
49.5
40.3
23.6
@@
w/o Weight Pred.
44.7
33.8
30.8
39.9
82.5
12.2
22.2 40.8
68.5
100.0 74.0
19.3
60.7
74.9
24.4
15.5
17.4
20.3
39.4
41.7
25.5
w/o SAM Integration
44.2
29.4
32.3
38.5
82.2
12.7
19.7
36.0
70.3
98.8
68.3
26.1 67.4 79.5
23.6
15.5
24.8
29.6
53.5
44.7
23.1
PartSTAD(Ours)
45.6 35.5 33.2
35.3
82.5
11.6
22.1
40.6
72.0
84.3
71.6
19.9
65.5
77.4
28.9
13.8
26.5 33.7
39.0
50.3 26.5
Table 2: Part-aware mAPs(%) of the instance segmentation results on the PartNet-Mobility dataset. Ours achieves 4.0%p improvement in mean part-aware mAPs compared to PartSLIP [32], with consistent enhancements across all categories. Please refer to the appendix for the complete table with results for all 45 categories.
Method
Mean
Storage Furniture
Table Faucet Chair
Trash Can
Switch Toilet
Eye- glasses
Globe Cart Window
Dis- penser
Bottle Laptop
Coffee Machine
USB Remote Pen
Dish- washer
Scissors
SAM3D [62]
12.1
1.2
10.6
24.0
5.5
9.6
5.4
3.7
27.6
3.9
8.6
18.0
20.7
16.0
1.5
3.6
22.6
1.0
8.0
6.7
7.2
PartSLIP [32]
38.9
29.7
28.7
42.8
80.7
22.6
21.2
35.4
71.8
92.1
69.5
23.6
59.0
73.4
19.5
9.8
20.5
19.9
40.8
39.0
27.0
@@
w/o Weight Pred.
42.6
35.9
29.2
37.1
81.1
24.5
22.2
40.0
72.1
100.0 74.0
19.3
59.9
74.9
26.5
9.5
15.2
20.3
35.8
41.9
29.4
w/o SAM Integration
42.6
34.8
27.6
37.9
81.2
24.4
19.7
36.6
72.5
98.8
68.3
26.1 69.0 79.5
27.4
9.4
27.3
29.6
47.2
47.1
24.9
PartSTAD (Ours)
44.1 41.7
28.2
35.6
83.3
22.8
22.4 41.0 74.0
84.3
71.2
19.5
65.5
76.2
34.1
9.4
23.5
33.7
34.5
49.7
26.2
Table 3: Part-agnostic mAPs(%) of the instance segmentation results on the PartNet-Mobility dataset. Ours demonstrates 5.2%p improvement in mean part-agnostic mAPs compared to PartSLIP [32], surpassing SAM3D with a mean mAP that is more than three times larger. Please refer to the appendix for the complete table with results for all 45 categories.
@@ 5.3 Instance Segmentation Results
Following the instance segmentation idea of PartSLIP, we also perform instance segmentation using our method and compare the quantitative results with those of other methods in Tab. 2 and Tab. 3, reporting part-aware mAPs and part-agnostic mAPs, respectively. Fig. 4 also displays qualitative comparisons across different methods. Please refer to the appendix for the complete table of all 45 categories. Similar to semantic segmentation, ours achieves 4.0%p and 5.2%p improvements for both average part-aware mAPs and part-agnostic mAPs, respectively. Both the weight prediction network and SAM mask integration also exhibit meaningful improvements in performance, as their influence is depicted in the ablation study results. Compared to SAM3D [62], whose part-agnostic mAPs are shown in Tab. 3, our method achieves much better performance with more than three times greater mAP. Qualitative results also illustrate the outstanding performance of our method, accurately identifying all the tiny instances of parts, such as keys on keyboards, buttons on remotes, and clear boundaries of these instances.
6 Conclusion

### v3

@@ Abstract
We introduce PartSTAD, a method designed for the task adaptation of 2D-to-3D segmentation lifting. Recent studies have highlighted the advantages of utilizing 2D segmentation models to achieve high-quality 3D segmentation through few-shot adaptation. However, previous approaches have focused on adapting 2D segmentation models for domain shift to rendered images and synthetic text descriptions, rather than optimizing the model specifically for 3D segmentation. Our proposed task adaptation method finetunes a 2D bounding box prediction model with an objective function for 3D segmentation. We introduce weights for 2D bounding boxes for adaptive merging and learn the weights using a small additional neural network. Additionally, we incorporate SAM, a foreground segmentation model on a bounding box, to improve the boundaries of 2D segments and consequently those of 3D segmentation. Our experiments on the PartNet-Mobility dataset show significant improvements with our task adaptation approach, achieving a 7.0%p increase in mIoU and a 5.2%p improvement in mAP50 for semantic and instance segmentation compared to the SotA few-shot 3D segmentation model. The code is available at https://github.com/KAIST-Visual-AI-Group/PartSTAD.
Keywords:
part segmentation few-shot learning 3D deep learning
@@ 3.1 PartSLIP [33]
Voting for Super Points.
@@ 4 PartSTAD: Task Adaptation for 2D-to-3D
In the supplementary material, we demonstrate that the choice of the loss function, as compared to other alternatives such as the cross-entropy loss, is crucial to achieving a substantial improvement in our task adaptation.
@@ 4.3 SAM [21] Mask Integration
Input Refer to caption
GT Refer to caption
SATR [2] Refer to caption
SATR [2]+SP Refer to caption
PartSLIP [33] Refer to caption
PartSTAD (Ours) Refer to caption
@@
Table 1: Quantitative comparison of the semantic segmentation results (mIoU(%)) on the PartNet-Mobility dataset. Ours achieves a 7.0%p improvement in average mIoU compared to PartSLIP [33], the SotA few-shot 3D segmentation method, while consistently increasing mIoU across all categories. Please refer to the supplementary material for the complete table with results for all 45 categories.
Method mIoU Storage Furniture Table Chair Switch Toilet Laptop USB Remote Scissors
SATR [2] 29.3 20.6 23.3 33.1 21.4 17.6 11.2 30.2 17.2 36.8
SATR [2]+SP 34.8 28.9 28.0 37.7 37.0 22.1 12.4 33.4 28.0 43.0
PartSLIP [33] 58.0 52.3 44.6 82.8 52.1 50.4 31.2 52.1 36.6 61.4
@@
w/o Weight Pred. 61.9 56.6 45.0 85.0 51.9 56.6 31.5 57.1 36.6 60.8
w/o SAM Integ. 62.1 54.0 45.7 83.1 53.0 49.4 33.6 53.6 46.8 67.5
PartSTAD (Ours) 65.0 59.5 47.8 85.3 57.9 57.5 34.6 59.9 53.4 68.5
@@ 5 Experiment Results
In the experiments, we compare our method with the SotA 2D-prior-based zero-shot/few-shot 3D segmentation models using the PartNet-Mobility [58] dataset and OmniObject3D [56] dataset. In the supplementary material, we present additional results on the ablation study, outcomes with scanned data, and comprehensive results covering all PartNet-Mobility categories.
@@ 5.1 Experiment Setup
Dataset.
We use the PartNet-Mobility [58] dataset, which is also used in the experiments of PartSLIP [33]. The PartNet-Mobility [58] dataset includes 45 object categories. For each category, the training split includes 8 shapes, while the test split has a range of the number of objects from 6 to 338, totaling 1,906 objects across all 45 categories. Following PartSLIP, 8 shapes from the test set of each category are designated as the validation set. To obtain the 2D bounding boxes for each object, we use 10 images from fixed viewpoints across all our experiments. Note that the PartNet-Mobility [58] dataset is identical to the subset of PartNet-Ensembled dataset used to train and evaluate PartSLIP [33]. (The remaining part of PartNet-Ensembled was only used to train other supervised baseline models, not PartSLIP [33].)
Evaluation Metrics.
@@
For instance segmentation, we compute two types of mAP: part-aware mAP and part-agnostic mAP. For part-aware mAP, similar to mIoU in semantic segmentation, we first calculate the mAP for each semantic part with instances having the part label. We then average them for each category, and then across categories. For part-agnostic mAP, we do not consider the semantic labels and directly compute the category-level mAP with part instances in the same object category, and then average them. All mAP values correspond to mAP50 with an IoU threshold of 50%.
Baselines.
@@
Ablation Study.
We also conduct an ablation study, comparing our method with two major components being ablated: the weight prediction (Sec. 4.2) and the SAM mask integration (Sec. 4.3). In the supplementary material, we also compare our method with cases using cross-entropy loss instead of mRIoU loss (Sec. 4.1) and using GLIP bounding box confidence score instead of our learned weights in the weighted voting (Sec. 4.2). Note that if both weight prediction and SAM mask integration are excluded, our PartSTAD is identical to PartSLIP [33].
@@ 5.2 Semantic Segmentation Results
Tab. 1 and Fig. 3 present the quantitative and qualitative comparisons of part semantic segmentation results, respectively. Please refer to the supplementary material for the complete table encompassing all 45 categories. In comparison to PartSLIP [33], a SotA few-shot 3D segmentation method, we achieve a significant 7.0%p mIoU improvement over the entire set of 45 categories. Furthermore, for each specific category, we consistently demonstrate improvement, surpassing by more than 15%p in some categories (e.g., Remote). Qualitatively, the results also highlight the effectiveness of our task adaptation and the SAM mask integration in segmenting parts more precisely, even for small (e.g., Camera, Chair) and thin (e.g., Clock) parts of objects.
@@
Input Refer to caption
GT Refer to caption
SAM3D [65] Refer to caption
PartSLIP [33] Refer to caption
PartSTAD (Ours) Refer to caption
@@
Table 2: Part-aware mAPs(%) of the instance segmentation results on the PartNet-Mobility dataset. Ours achieves 4.0%p improvement in mean part-aware mAPs compared to PartSLIP [33], with consistent enhancements across all categories. Please refer to the supplementary material for the complete table with results for all 45 categories.
Method mAP Storage Furniture Table Chair Switch Toilet Laptop USB Remote Scissors
PartSLIP [33] 41.6 29.1 32.6 82.2 21.2 36.2 17.8 20.9 19.9 23.6
@@
w/o Weight Pred. 44.7 33.8 30.8 82.5 22.2 40.8 24.4 17.4 20.3 25.5
w/o SAM Integ. 44.2 29.4 32.3 82.2 19.7 36.0 23.6 24.8 29.6 23.1
PartSTAD(Ours) 45.6 35.5 33.2 82.5 22.1 40.6 28.9 26.5 33.7 26.5
Table 3: Part-agnostic mAPs(%) of the instance segmentation results on the PartNet-Mobility dataset. Ours demonstrates 5.2%p improvement in mean part-agnostic mAPs compared to PartSLIP [33], surpassing SAM3D with a mean mAP that is more than three times larger. Please refer to the supplementary material for the complete table with results for all 45 categories. (*SAM3D is a zero-shot method.)
Method mAP Storage Furniture Table Chair Switch Toilet Laptop USB Remote Scissors
SAM3D* [65] 12.1 1.2 10.6 5.5 5.4 3.7 1.5 22.6 1.0 7.2
PartSLIP [33] 38.9 29.7 28.7 80.7 21.2 35.4 19.5 20.5 19.9 27.0
@@
w/o Weight Pred. 42.6 35.9 29.2 81.1 22.2 40.0 26.5 15.2 20.3 29.4
w/o SAM Integ. 42.6 34.8 27.6 81.2 19.7 36.6 27.4 27.3 29.6 24.9
PartSTAD (Ours) 44.1 41.7 28.2 83.3 22.4 41.0 34.1 23.5 33.7 26.2
@@ 5.3 Instance Segmentation Results
Following the instance segmentation idea of PartSLIP, we also perform instance segmentation using our method and compare the quantitative results with those of other methods in Tab. 2 and Tab. 3, reporting part-aware mAPs and part-agnostic mAPs, respectively. Fig. 4 also displays qualitative comparisons across different methods. Please refer to the supplementary material for the complete table of all 45 categories. Similar to semantic segmentation, ours achieves 4.0%p and 5.2%p improvements for both average part-aware mAPs and part-agnostic mAPs, respectively. Both the weight prediction network and SAM mask integration also exhibit meaningful improvements in performance, as their influence is depicted in the ablation study results. Compared to SAM3D [65], whose part-agnostic mAPs are shown in Tab. 3, our method achieves much better performance with more than three times greater mAP. Qualitative results also illustrate the outstanding performance of our method, accurately identifying all the tiny instances of parts, such as keys on keyboards, buttons on remotes, and clear boundaries of these instances.
Figure 5: Qualitative comparison of semantic segmentation results on OmniObect3D [56] dataset, a high quality real scanned 3D objects dataset. Our PartSTAD predicts precise boundaries, even in the case that the object has an appearance significantly different from training data such as box category.
5.4 Part Segmentation with Real Scanned Dataset
Fig. 5 displays additional results on the OmniObject3D [56] dataset, a high-quality dataset featuring real-scanned 3D objects. We present qualitative results only, as ground truth segmentation for OmniObject3D [56] is unavailable. Note that our PartSTAD accurately identifies parts in real scans, demonstrating its effectiveness beyond synthetic data. Particularly impressive is its performance in the box category depicted in Fig. 5, where despite significant discrepancies in appearance compared to the training data, PartSTAD achieves remarkably accurate predictions.
6 Conclusion and Future Work